OpenStudy (anonymous):

What is the multiplicative inverse of 2 1/2?

zepdrix (zepdrix):

Hey Audie :) The inverse of multiplication is division. Example: The inverse of \(\large\rm \times5\) multiplying something by 5 is \(\large\rm \div5\) dividing by 5. Division is the same thing as `multiplying by the reciprocal`. That's the way we would prefer to think of it. So we would say that the multiplicative inverse of \(\large\rm 5\) is \(\large\rm \frac{1}{5}\).

zepdrix (zepdrix):

So to find the multiplicative inverse, I had to "flip" the fraction. Oh I guess we would have to think of 5 as \(\large\rm \frac{5}{1}\) in order to flip it. For your specific problem, you'll need to write your mixed number as an improper fraction before you can flip it.

zepdrix (zepdrix):

\[\large\rm 2\frac{1}{2}=\frac{5}{2}\]And then flip it, ya? Any confusion there? :O